A hiker walks from the drop-off and initially covers 3000 m, 35° north of west. He then traveled another 5500 m heading to 10° north of west. Finally, he reached his destination by walking 1500 m due north. What is his total displacement?

Scale: 1 cm = 1 000 m



1
Expert's answer
2021-08-31T17:33:11-0400
"s_x=3000\cos{35}+5500\cos{10}=7873.9\ m\\
s_y=3000\sin{35}+5500\sin{10}+1500=4175.8\ m\\s=\sqrt{7873.9^2+4175.8^2}=8913\ m\\
\theta=\arctan{\frac{4175.8}{7873.9}}=28\degree \text{ N of W}"